OpenStudy (anonymous):

Factor the polynomial completely. 2a^2 - 162 Show steps please! :)

OpenStudy (anonymous):

2(a^2-81) =2(a^2-9^2) =2(a+9)(a-9)

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Remember: \[a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

The sum of squares.. Ugh, so many formulas. Lol, thank you! :)

OpenStudy (compassionate):

2(a^2 - 81) = 2(a + 9)(a + 9) 162 is positive, adding a negative would reduce it.

OpenStudy (anonymous):

No, it is subtraction of squares... He he he...

OpenStudy (anonymous):

Difference of squares* Thanks @Compassionate :)

OpenStudy (anonymous):

Haha, yeah.. I just caught myself.. :\

OpenStudy (anonymous):

It happens sometimes...

OpenStudy (anonymous):

@Compassionate i think waterneyes is rigth

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/pdf/Algebra_Cheat_Sheet_Reduced.pdf @LynnMichael

OpenStudy (anonymous):

Oh, sweet! Thank you so much @Outkast3r09 !! :)

OpenStudy (anonymous):

no problem